OpenStudy (anonymous):

what is the simplified form of (3.25x10^3)(7.8x10^6) written in scientific notation?

OpenStudy (anonymous):

these are the answers a) 2.535x10^10 b)2.535x10^18 c)2.535x10^19 d)2.535x10^9

OpenStudy (anonymous):

first multiply your front terms: 3.25*7.8 = 25.35 however, in scientific notation, you want your largest term to be in the one's place to simplify your number. This can be rewritten as 25.35 = 2.535*10 or 2.535x10^1 Now all you can to do is add the exponents of all the 10 terms You have 10^3, 10^6, and now 10^1 3+6+1 = 10 That means your final answer is 2.525*10^10

OpenStudy (anonymous):

so the answer would be a ???

OpenStudy (anonymous):

it would appear so